OpenStudy (anonymous):

Your English class has 27 students. You would like to treat the class to cookies. You would like each person to have 3 cookies. Which expression would help you figure out how many cookies you would need to bake?

OpenStudy (nubeer):

let number of students =x number of cookies =y eact student will get 3 cookies.. so y =3x. i suppose this should be the answer

OpenStudy (accessdenied):

I agree with the above answer, although I think it makes more sense when observing units. We have a number of students (x), a number of cookies (y), and a rate (3): cookies/student. \( \displaystyle \frac{\text{cookies}}{\cancel{\text{student}}} \times \cancel{\text{students}} = \text{cookies}\) \(\displaystyle \frac{3 \text{ cookies}}{1 \text{ student}} \times x \text{ students} = y \text{ cookies}\) \(3x = y\) I am just offering a potentially useful way of looking at these problems. I found this more intuitive recently in other Algebra problems. :)

OpenStudy (nubeer):

lol so was i wrong or something.. or not the right way to explain?

OpenStudy (accessdenied):

Nah, you are correct. I just wanted to offer a second way to get the answer that I like / I think is really interesting. :)

OpenStudy (nubeer):

ohh lol i use it mostly in physics and thermodynamics.. lol its more like my trick :P